# Semidiscrete optimal transport with unknown costs

### v3

@@ 6.2 Geographical partitioning
For each $k=1,...,K$, let $x_{k}\in\left[0,1\right]^{2}$ be a fixed location (“facility”) in the unit square. A user appears at some random location $X\in\left[0,1\right]^{2}$ and is assigned to a facility. The cost incurred by assigning the user to facility $k$ is $c_{k}\left(X\right)=\|X-x_{k}\|_{2}$, the Euclidean distance between the facility and the user. The goal is to choose an assignment rule that minimizes the expected distance subject to the probabilistic targets (which, in this context, require each facility to serve a certain proportion of the population). It is well-known in computational geometry that the optimal policy can be visualized as an additively weighted Voronoi diagram (Carlsson et al., 2016) whose weights are the values $g^{*}_{k}$. The diagram partitions the unit square into regions, each containing a single facility, to which all users appearing in that region are assigned.
@@
We considered two instances of this model where $\varepsilon\sim\mathcal{N}\left(0,0.02^{2}\right)$ and $X$ is uniformly distributed on $\left[0,1\right]^{2}$. Note that, even with the normality assumption, the residual error in (37) is no longer sub-Gaussian. The instances used different facility locations $x_{k}$ and targets $p$. The settings of the semi-myopic policy were the same as in Section 6.1.

### v4

@@ 6.2 Geographical partitioning
For each k=1,,Kk=1,...,K, let $x_{k}\in\left[0,1\right]^{2}$ be a fixed location (“facility”) in the unit square. A user appears at some random location $X\in\left[0,1\right]^{2}$ and is assigned to a facility. The cost incurred by assigning the user to facility $k$ is $c_{k}\left(X\right)=\|X-x_{k}\|_{2}$, the Euclidean distance between the facility and the user. The goal is to choose an assignment rule that minimizes the expected distance subject to the probabilistic targets (which, in this context, require each facility to serve a certain proportion of the population). It is well-known in computational geometry that the optimal policy can be visualized as an additively weighted Voronoi diagram (Carlsson et al., 2016) whose weights are the values $g^{*}_{k}$. The diagram partitions the unit square into regions, each containing a single facility, to which all users appearing in that region are assigned.
@@
We considered two instances of this model where $\varepsilon\sim\mathcal{N}\left(0,0.02^{2}\right)$ and $X$ is uniformly distributed on $\left[0,1\right]^{2}$. Note that, even with the normality assumption, the residual error in (46) is not sub-Gaussian. The instances used different facility locations $x_{k}$ and targets $p$. The settings of the semi-myopic policy were the same as in Section 6.1.